in 1990, the population of the the United States was about two hundred fifty million people (250,000,000). The first release from census 2000 shows that the U.S. population is 281,400,000. What is the rate of change in the population during that 10 year period.
Answer #1

Population in 1990 = 250,000,000

Population in 2000 = 281,400,000

Time = 10 years

Let the rate of change of population in United States be i.

281400000 = 250000000 * (1+i)10

= 2814/2500 = (1+i)10

= 1.0119 = 1+i

= i = 0.0119.

So, the rate of change of population was 1.19% per year.

IF YOU FIND THIS ANSWER HELPFUL, PLEASE UP VOTE.